OpenStudy (anonymous):

Boyce invested an average of $450 per month since age 34 in various securities for his retirement savings. His investments averaged a 4% annual rate of return until he retired at age 63. Given the same monthly investment and rate of return, how much more would Boyce have in his retirement savings had he started investing at age 20? Assume monthly compounding. $113,320.00 $321,946.93 $78,624.00 $224,749.27

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

use the same formula, but u need to calculate it two times... and subtract

ganeshie8 (ganeshie8):

\(\large FV = C\bullet \dfrac{(1+i)^{nt}-1}{i}\)

ganeshie8 (ganeshie8):

first use it for t = 63-34 next use it for t = 63-20